OpenStudy (mayankdevnani):

Plz...!Plz...!Plz...!Plz...!Plz...!Plz...!Plz...!Plz...!Plz...!Plz...!Plz...!Plz...!Plz...!Plz...!Plz...!Plz...!Plz...!Plz...!Plz...!....... Someone HELP me!!!! for finding the solution for this question which is listed below: Q-) Find the HCF of 81 and 237 and express it as a linear combination of 81 and 237.

mathslover (mathslover):

HCF of 81 and 237 81 = 3 * 3 * 3 * 3 237 = 3 * 79 HCF = ?

OpenStudy (mayankdevnani):

i know the HCF of 81 and 237... which is listed below:(Using Euclid's Lemma) \[237=81\times2+75\] \[81=75\times1+6\] \[75=6\times12+3\] \[6=3\times2+0\] am i right????

mathslover (mathslover):

Yes! now 3 can be expressed as linear combination of 81 and 237 as 3 = 81x + 237y, where x and y are not unique.

OpenStudy (mayankdevnani):

hence, HCF is 3

mathslover (mathslover):

Got it mayank?

OpenStudy (anonymous):

HCF=3

OpenStudy (mayankdevnani):

got it!!! @mathslover and @KingAditya i mentioned it earlier!!!

mathslover (mathslover):

We have to use the HCF to represent it in the linear combination of 81 and 237 : So we can write the HCF as : 3 = 81 x + 237 y. Lemme explain more.

OpenStudy (mayankdevnani):

@mathslover is this final answer????

Parth (parthkohli):

Err, it's more like the whole algorithm.

OpenStudy (mayankdevnani):

i don't understand!!!! @Eyad plz explain it by one by one!!! plzz///

OpenStudy (anonymous):

Which step you don't understand ?

OpenStudy (mayankdevnani):

after finding the HCF of 81 and 273 .......then then this whole process which is below than hcf of 81 and 273 i don't understand

mathslover (mathslover):

\(\begin{array}{} 237 = 81 × 2 + 75 \\ 81 = 75 × 1 + 6 \\ 75 = 6 × 12 + 3\\ 6 = 3 × 2 + 0 \end{array}\) Now it is clear that HCF = 3 In the step 3 : \(\begin{array}{} 75 = 6 \times 12 + 3 --- (\text{step 5}) \\ \text{and from step 2: } \space \\ 81 = 75 \times 1 + 6 \\ \text{thus we can write step 5 as : } \\ 3 = 75 – (81 – 75 × 1) × 12\\ \implies 3 = 75 – (81× 12 – 75 × 12)\\ \implies 3 = 75 × 13 – 81× 12 ... \space (\text{step 6})\\ \text{From step 1:}\\ 75 = 237 - 81 \times 2 \\ \text{thus from step 6 :} \\ 3 = (237 – 81 × 2) × 13 – 81× 12\\ \implies 3 = (237 × 13 – 81 × 26) – 81× 12 \\ \implies 3 = 237 × 13 – 81 × 38 \\ \implies \text{H.C.F of 237 and 81 = } 237 \times 13 + 81 \times (-38) \\ \end{array}\) ∴HCF, 3 can be expressed as linear combination of 81 and 237 as 3 = 81x + 237y, where x and y are not unique.

mathslover (mathslover):

I hope this clears your doubt mayank.

mathslover (mathslover):

Thanks to meritnation too.

OpenStudy (anonymous):

This whole process was proof to your query , 81 = 3 x 27 = 3 x 9 x 3 = 3 x 3 x 3 x 3 = 34 Again, 237 = 79 x 3 = 3 x 79 Hence, H.C.F = 3 ,therefore 81*3/2-273*1/2 =243/2-273/2 =6/2 =3

mathslover (mathslover):

I think @mayankdevnani is getting problem in understanding what is linear combination and what does the second part mean? Is it so Mayank?

OpenStudy (mayankdevnani):

yaaa.. this is true friend.....nature

mathslover (mathslover):

just like : 3 = 2x + 3y, similarly in this question we will represent HCF (=3) in the form of : 81x + 237 y

OpenStudy (mayankdevnani):

plz wait!!!! i need time to understand!!!! plzzz wait

mathslover (mathslover):

Take your time my friend.

OpenStudy (mayankdevnani):

3=(81-75*1)*12.....where this eq. comes?????

OpenStudy (mayankdevnani):

@mathslover plz expalin it one by one with your comment and then my comment!!!!! can you???

mathslover (mathslover):

yes! Of course! Ready to help anyone anytime. Linear combination , let us first understand this term .

OpenStudy (mayankdevnani):

yaaa.plzzz

mathslover (mathslover):

ax + by = c [ just like linear equation ]

OpenStudy (mayankdevnani):

okkkk

OpenStudy (anonymous):

Can i join too.. @mathslover ?

mathslover (mathslover):

See now we want to write HCF = 3 = c as in terms of ax + by. we have a = 81 and b = 237 so : 3 = 81 x + 237 y , right?

OpenStudy (mayankdevnani):

right

mathslover (mathslover):

So now we want to find x and y, scroll up mayank and check the latex work I presented. It is clear that: H.C.F. of 237 and 81 = 237 × 13 + 81 × (–38) right?

mathslover (mathslover):

So we have : 3 = 237 * 13 + 81 * (-38) as the answer for the second part.

OpenStudy (mayankdevnani):

pzz wait!!!

mathslover (mathslover):

I have to go now, sorry friend! May be @KingAditya can help you if you get any problem in understanding this. Hope you understand my problem :( Thanks!

OpenStudy (mayankdevnani):

i don't understand yet!!!

OpenStudy (anonymous):

I can make it clear to u..

OpenStudy (mayankdevnani):

plzzzz

OpenStudy (anonymous):

U got what @mathslover told u?

OpenStudy (mayankdevnani):

237 × 13 + 81 × (–38)...where this comes...@mathslover

OpenStudy (mayankdevnani):

@mathslover

OpenStudy (anonymous):

It is the 2nd part of the answer...to show the linear combination

OpenStudy (mayankdevnani):

@mathslover i understand!!!! plz go on!!!! but one thing arise:- where 13 comes in eq. scroll up (STEP-6)

OpenStudy (mayankdevnani):

@mathslover r u there????????? i need help!!! my friend

OpenStudy (mayankdevnani):

@mathslover plz tell

OpenStudy (chihiroasleaf):

are you asking the step 6?

OpenStudy (mayankdevnani):

yaa...correct @chihiroasleaf

OpenStudy (mayankdevnani):

@Eyad where does 13 comes in step 6?????thats was my main problem

OpenStudy (mayankdevnani):

@mukushla and @gerryliyana plz help me!!!!

OpenStudy (mayankdevnani):

@ParthKohli and @hartnn plzz help me!!!!

OpenStudy (mayankdevnani):

@mathslover @hba @ParthKohli @Eyad @electrokid

OpenStudy (mayankdevnani):

@UnkleRhaukus @ParthKohli

Parth (parthkohli):

I guess it's more of mathslover's thing. I haven't used the Extended Euclidean Algorithm a lot.

OpenStudy (mayankdevnani):

oh!!!! kkk

OpenStudy (mayankdevnani):

can you help me @mathslover and @mukushla ...little help!!!! not so big help

mathslover (mathslover):

Sorry, cant help. Busy right now

OpenStudy (mayankdevnani):

in which work?????

OpenStudy (mayankdevnani):

@dmezzullo

OpenStudy (mayankdevnani):

@amistre64

OpenStudy (amistre64):

this is an application of a theorem in number theory; the gcd of 2 numbers can be expressed as a linear combination of those numbers.

OpenStudy (mayankdevnani):

yaa!!!

OpenStudy (amistre64):

spose gcd(a,b) = c then ax+by = c with x,y having infintie solutions

OpenStudy (mayankdevnani):

okkk

OpenStudy (mayankdevnani):

@amistre64 will you explain me mathslover sol.???

OpenStudy (chihiroasleaf):

you're still asking the step 6, right?

OpenStudy (mayankdevnani):

right!!!

mathslover (mathslover):

@mayankdevnani

OpenStudy (mayankdevnani):

what????

mathslover (mathslover):

I hope it helps you mayank.

OpenStudy (mayankdevnani):

i can't open it!!!!

mathslover (mathslover):

You have to open it with : internet explorer.

OpenStudy (mayankdevnani):

it contains what????

OpenStudy (chihiroasleaf):

it's just simple distribution and multiplication rule \[3=75–(81×12–75×12)\] \[3 = 75 - (81 x 12) + (75 x 12)\] \[3=75 + 75 \times 12 –81\times12 \] \[3 = 75 (1+12) - 81 \times 12\] \[3 = 73 \times 13 - 81 \times 12 \] do you get it? :)

mathslover (mathslover):

Right click on it : Save link as , save it. Then open it with internet explorer.

mathslover (mathslover):

It contains your solution Mayank

OpenStudy (mayankdevnani):

nice hand writing man!!!!

OpenStudy (mayankdevnani):

plz wait for a min.

mathslover (mathslover):

3=75–(81×12–75×12) 3 = 75 - (81 * 12) + (75 * 12) 3 = 75 + 75 * 12 - (81 *12) 3 = 75(12+1) - (81*12) 3 = 75 * 13 - 81 * 12 this is the explaination for step 6.

OpenStudy (amistre64):

im not to good at trying to explain other peoples ideas :) but if we do a euclidian run on this we get 237 = 81(2) + 75 81 = 75(1) + 6 75 = 6(12) + 3 <-- gcd=3 6 = 3(2) + 0 81x + 237y = 3 to find the smaller of 81 and 237, a specific value for x, we can run the reverse euclidean algorithm 0 0 1 0 - 2(1) = -2 1 -1(-2)= 3 -2-12(3) = -38 <-- x' = -38; this leaves is with y' = 13 as one soluton

OpenStudy (mayankdevnani):

thnx... @mathslover but in your work(file) that you upload , i can't understand your last 2 line!!!! i mean how did you get 3=237*13-81*38

OpenStudy (mayankdevnani):

@amistre64 and @chihiroasleaf thank you.....

OpenStudy (mayankdevnani):

@amistre64 and @chihiroasleaf can you help me????

mathslover (mathslover):

I have : 3 = (237 * 13 - 81 *26 ) - 81 * 12 3 = 237 * 13 - 81 * 26 - 81 * 12 3 = 237 * 13 - 81*(26+12) 3 = 237 * 13 - 81 * (38) Got it Mayank?

OpenStudy (mayankdevnani):

i got it @mathslover

OpenStudy (mayankdevnani):

i got it @mathslover

OpenStudy (amistre64):

i believe there is a setup such that:\[x=x'+\frac bcn\]\[y=y'-\frac acn\] but thats prolly incorrect since its from memory, but it looks something like that

OpenStudy (mayankdevnani):

thnx.... @amistre64 , @chihiroasleaf , @Eyad and special thanks to my friend.. @mathslover

mathslover (mathslover):

Amistre, I liked your method also. Welcome mayank.

OpenStudy (mayankdevnani):

can i ask you last question...its so simple!!!! mathslover

OpenStudy (mayankdevnani):

@mathslover

mathslover (mathslover):

Hmn, Yes ask mayank...

OpenStudy (amistre64):

i believe there is a setup such that:\[x=ax'+\frac bcn\] but thats prolly incorrect since its from memory, but it looks something like that